OpenStudy (anonymous):

write an equation leaving answer in slope form for a slope of -2/3, y intercept (0,5)

OpenStudy (nubeer):

y-y1 = m(x-x1) m =slope , x1 = 0 , y1 = 5 just plug in the values u will get the equation